OpenStudy (anonymous):

On the grid, the distance between the points located at (-3,-4) and (-8,1) is roughly equal to ___. A) 6.21 B) 5.83 C) 4 D) 7.07 @satellite73

OpenStudy (anonymous):

\[\sqrt{(-3+8)^2+(-4-1)^2}\] and a calculator

OpenStudy (anonymous):

I got 25 - 25, but i haven't square rooted it yet

OpenStudy (anonymous):

-3 + 8 = 5 * 5 -4-1 = -5 * 5 25 - 25

OpenStudy (anonymous):

that is a plus sing in there, not a minus sign \[\sqrt{50}\]

OpenStudy (anonymous):

7.07?

OpenStudy (anonymous):

looks reasonable

OpenStudy (anonymous):

Yep! I've only got a few more :p

OpenStudy (anonymous):

Thanks a lot satellite (: